### v1

# Failure Makes the Agent Stronger: Enhancing Accuracy through Structured Reflection for Reliable Tool Interactions

Abstract
Tool-augmented large language models (LLMs) are typically trained via supervised imitation learning or coarse-grained reinforcement learning, approaches that primarily optimize one-shot tool calls. Existing practices of self-reflection largely rely on heuristic prompting or unidirectional reasoning traces: the model is encouraged to “think more,” rather than to treat error diagnosis and correction as a learnable capability. This makes them fragile in multi-turn interaction settings—once a call fails, the model tends to repeat the same mistake instead of recovering. To address this issue, we propose structured reflection, which transforms the “from error to repair” process into a first-class, controllable, and trainable action. The agent produces a concise yet precise reflection process: specifically, the model diagnoses the error based on evidence from the previous step and then proposes a correct and executable follow-up call. During training, we combine DAPO and GSPO’s objective functions and design a more principled reward mechanism tailored to tool calling, optimizing the stepwise strategy Reflect $\to$ Call $\to$ Final. To evaluate this capability, we introduce Tool-Reflection-Bench, a lightweight benchmark dataset that programmatically verifies structural validity, executability, parameter correctness, and result consistency. Tasks in the benchmark are constructed as miniature trajectories of Erroneous Call $\to$ Reflection $\to$ Corrected Call and are split into disjoint training and testing sets. Experiments on BFCL v3 and Tool-Reflection-Bench show that our method achieves significant improvements in multi-turn tool-call success rates and error recovery, while also reducing redundant calls. These results demonstrate that making reflection explicit and treating it as an optimization objective can substantially enhance the reliability of tool interaction, providing a reproducible pathway for agents to grow stronger by learning from failure.
footnotetext: * Equal Contribution. ✉Corresponding Authors. $\dagger$} Project Leader.
1 Introduction
The integration of external tools with large language models through tool calling represents a significant breakthrough in the development of agents. It transforms large language models from mere text generators into highly practical tools for interacting with humans [1, 2], significantly enhancing the ability of AI agents to solve complex real-world tasks [3, 4, 5]. Tool calling bridges the gap between the vast internal knowledge of LLMs and external resources, enabling LLMs to access up-to-date information, perform delicate computations, and more, thereby unlocking their broad potential for applications across multiple domains [6, 7, 8].
Currently, the training of tool-call capabilities in large language models typically relies on supervised fine-tuning and reinforcement learning [9, 10], where these methods optimize the ability for single-turn tool calls through carefully designed reward mechanisms. However, these approaches face several challenges in the context of tool calling. First, the issue of rewards in tool calling is particularly prominent—small errors in parameter selection or formatting often render the entire function call invalid, thus limiting the effective learning signal [11]. Second, existing methods generally rely on unidirectional reasoning, which, while sufficient for simpler scenarios, has clear limitations: when LLMs make mistakes during tool calls, they often struggle to locate the root cause of the error [12]. While generating correct function calls is crucial, it is even more important for LLMs to learn how to identify and correct their own mistakes [13].
To address the above-mentioned issues, we propose an innovative reflection process aimed at error localization and correction through explicit reflection steps, which differs from existing forward reasoning methods. Specifically, we design a process in which the LLM intentionally makes mistakes during tool calls, carefully crafts reflection content based on the errors, and then generates the correct call. Through this approach, we transform the self-correction ability of large models from a heuristic process [14] into a clear, trainable capability. Our training approach is primarily reinforcement learning–based. During the reinforcement learning process, we specifically design a customized reward mechanism tailored for tool-calling scenarios, with a particular emphasis on multi-turn interactions. Concretely, the reward design encompasses multiple dimensions, including format reward, tool-name reward, parameter reward, and semantic reward of reflection, which together provide the model with multi-dimensional feedback and effectively guide its learning, and we further combine DAPO’s decoupled clipping range and dynamic sampling—expanding exploration while skipping near-zero-advantage rollouts—with GSPO’s sequence-level importance sampling and same-granularity clipping, which avoids token/sequence mismatch and stabilizes optimization. With this training methodology, our approach equips LLMs with genuine self-reflection and error-correction capabilities. On the BFCL v3 benchmark, our method yields significant improvements in LLM accuracy for multi-turn tool calling, thereby demonstrating its effectiveness in real-world applications.
We construct a Tool-Reflection-Bench based on the BUTTON dataset [15] style. First, we collected tool-call failure cases from real-world scenarios and various benchmarks, analyzing and summarizing several common failure patterns. Next, We selected several existing tool-call datasets [4, 16] and randomly combined them according to the call style of the BUTTON dataset and introduced these failure patterns into the data, disrupting the originally correct call processes to generate failure cases. Finally, we meticulously designed a reflection process to repair these failures, resulting in successful tool calls. The training set includes the complete process described above to train LLMs to achieve true self-correction capabilities, while the test set only contains the first two steps, used to evaluate the self-correction abilities of the LLMs. By constructing the Tool-Reflection-Bench in this manner, combined with our custom reward mechanism for tool calling, we have made breakthroughs in LLMs’ self-correction abilities during training. Particularly in multi-turn tool-calling scenarios, we observed significant improvements in accuracy. Through the reasoning process from failure to correction, LLMs can more effectively identify and learn from potential mistakes, thus enhancing the model’s stability and robustness in interactions. This makes the agent’s behavior more robust and powerful.
In summary, our contributions are as follows:
  • We introduce an explicit, trainable reflection process that diagnoses the cause of a failed tool call using prior evidence and proposes a corrected, executable call. This transforms the "from failure to repair" process from a heuristic method into a learnable action strategy, enabling LLMs to genuinely possess self-reflection and error-correction capabilities, thereby enhancing the agent’s multi-turn interactions with users.
  • We design a more effective reward mechanism for tool call, tailored for RL training, using a GRPO-style objective function. This approach employs multi-dimensional rewards—format executability, tool name accuracy, parameter correctness, and semantic consistency—to mitigate sparse rewards and propagate signals across multi-turn trajectories.
  • We propose Tool-Reflection-Bench, which collects failure patterns from real interaction scenarios and benchmark datasets, injects perturbations into correct calls, and attaches a reflection process to repair the calls. This allows for training LLMs in their Self-Correction ability in tool-calling scenarios.
  • Our method significantly improves the accuracy of multi-turn tool calls and the ability to recover from tool call errors, while maintaining competitive single-turn tool call performance. We validate this by experiments on BFCL v3 [17] and Tool-Reflection-Bench.
2 Related Works
2.1 Tool-augmented Large Language Models
Integrating external tools into large language models has become a key approach to enhancing their functionality, surpassing the simple task of text generation. Traditional LLMs are limited by static knowledge, constrained to the data they were trained on. However, tool-augmented models extend the capabilities of LLMs by enabling them to interact with external resources [18, 19] (such as APIs [20], databases, and computational engines) through tool calls. This extension allows LLMs to access real-time data, perform external computations, and even interface with external hardware, making them more practical for solving complex real-world tasks that require dynamic information or specific external operations [21]. ToolBench [4] demonstrates the feasibility of integrating external tool calls into LLMs. Through such systems, LLMs can handle more specialized tasks. However, one major challenge of tool augmentation is how to effectively train LLMs to use these tools. Existing training methods, such as supervised fine-tuning and reinforcement learning, typically focus on optimizing single tool calls. This type of interaction often does not involve multi-turn tool calls or responses, which makes the limitations of current methods particularly apparent when errors occur during tool usage. In such cases, the model’s ability to recover from errors becomes crucial.
2.2 Self-Correction in LLMs
Self-correction in large language models refers to the model’s ability to diagnose its own errors and correct them based on previous actions [22, 23]. However, this area has not been fully explored. Existing self-correction techniques mostly rely on heuristic methods or unidirectional reasoning processes [24].
Self-Refine framework [25], which involves having LLMs provide an initial response, followed by a reflection process where the model identifies flaws and makes improvements. Specifically, the same LLM acts as both the responder and the evaluator: the model first generates an initial response, then self-reflects and iteratively revises the output. This approach has been shown to enhance the performance of LLMs in certain domains. However, subsequent studies [26, 27] have found that relying solely on the model itself often fails to detect subtle errors. Some research [28, 29] has introduced auxiliary verifiers (such as additional models or mechanisms [30, 31]) to help check the correctness of the initial response. This external self-checking assistance avoids unnecessary repeated revisions, improving efficiency and enhancing the model’s reasoning and verification capabilities. However, this approach remains highly sensitive to the specific phrasing of the prompts, with different prompt wordings leading to varying results [23].
However, even though these methods have somewhat improved the capabilities of LLMs, their essence still relies on external feedback to assist in correcting their own errors. In contrast, our work introduces a reflection method that transforms self-correction into a trainable and controllable capability. The reflection process is an explicit part of the task, where the model actively evaluates its previous actions, identifies errors, and generates explicit corrections. This process is achieved through error localization, diagnosis, and correction, which goes beyond simple unidirectional reasoning and can be integrated into the training process. By providing supervisory signals during training, our approach enables LLMs to truly possess self-correction capabilities, which are then reflected in tool calling tasks.
3 Method
3.1 Tool-Reflection-Bench
The construction of Tool-Reflection-Bench consists of the following steps: perturbation-based disruptions, positive samples transformations, and the reflection repair process. The original positive samples are derived from BUTTON [15] transformations and self-constructed based on few-shot prompts. The entire benchmark is divided into a training set and a test set, with approximately 5,000 samples in the training set, in addition to the reflection-augmented data constructed as described above, the training set also contains a very small portion of original data drawn from BUTTON [15] and XLAM [18]. And around 1,000 samples in the test set, the test set is exclusively composed of perturbation-derived items and does not include raw, unperturbed positives from BUTTON or XLAM.
3.1.1 Perturbation-based Disruptions
Let the initial correct message sequence be
$D^{+}=\Big(m^{\mathrm{sys}}_{0},\,m^{\mathrm{usr}}_{1},\,m^{\mathrm{ast}}_{2},\,m^{\mathrm{tool}}_{3},\,m^{\mathrm{ast}}_{4},\,m^{\mathrm{tool}}_{5},\,\dots,\,m^{\mathrm{ast}}_{2k},\,m^{\mathrm{tool}}_{2k+1},\,\dots,\,m^{\mathrm{final}}_{n}\Big),$ (1)
where $m^{\mathrm{sys}}_{0}$ is the system prompt, $m^{\mathrm{usr}}_{1}$ the user query, $m^{\mathrm{ast}}_{2i}$ the assistant’s $i$-th tool call in structured form (e.g., <call>[{…},{…},...]</call>), $m^{\mathrm{tool}}_{2i+1}$ the tool return (JSON), and $m^{\mathrm{final}}_{n}$ the final answer.
We define a set of disruption operators
$\mathcal{P}=\{P_{1},P_{2},P_{3},P_{4}\},$ (2)
each operating on an assistant call $m^{\mathrm{ast}}_{2k}$ and instantiating a common failure mode:
  1. 1.
    $P_{1}$ call-order swap: replace the current tool call with the next-round tool call dialogue and force an error.
  2. 2.
    $P_{2}$ redundant call: repeat the same tool at the step (unchanged/irrelevant arguments) and force an error.
  3. 3.
    $P_{3}$ missing call: replace the intended tool by another tool and force an error.
  4. 4.
    $P_{4}$ argument error: randomly corrupt the arguments of a call (missing/typed/alias/boundary) and force an error.
These operators specify how a correct tool call can be broken.
3.1.2 Positive Samples Transformations
Given a clean trajectory $D^{+}$ and a chosen operator $P_{j}\in\mathcal{P}$ acting on step $2k$, we produce the negative (erroneous) context; no repair is performed in this step. We construct the erroneous call
$\tilde{m}^{\mathrm{ast}}_{2k}=\mathrm{ApplyPerturbation}\!\left(m^{\mathrm{ast}}_{2k},\,P_{j}\right),$ (3)
and simulate the tool’s error feedback with a LLM $\mathcal{L}$:
$\tilde{m}^{\mathrm{tool}}_{2k+1}=\mathcal{L}\!\left(\tilde{m}^{\mathrm{ast}}_{2k};\,\mathcal{L}\right).$ (4)
This yields the negative trajectory prefix
$D^{-}=\mathrm{Perturb}\!\left(D^{+},P_{j}\right)=\Big(m^{\mathrm{sys}}_{0},\,m^{\mathrm{usr}}_{1},\,\ldots,\,\tilde{m}^{\mathrm{ast}}_{2k},\,\tilde{m}^{\mathrm{tool}}_{2k+1}\Big),$ (5)
which will later serve as evidence of failure. At this stage, the item consists only of the broken call and its error signal.
3.1.3 Reflection Repair Process
Given a clean trajectory $D^{+}$ and its perturbed prefix $D^{-}$, we present the LLM with a paired view of the step-$2k$ evidence:
$\text{clean: }(m^{\mathrm{ast}}_{2k},\,m^{\mathrm{tool}}_{2k+1})\quad\text{vs.}\quad\text{broken: }(\tilde{m}^{\mathrm{ast}}_{2k},\,\tilde{m}^{\mathrm{tool}}_{2k+1}).$ (6)
The model outputs a response.
$\langle\texttt{reflect}\rangle r\langle/\texttt{reflect}\rangle,$ (7)
where $r$ briefly diagnoses the discrepancy, and $c$ proposes the fixed tool call. We then apply human supervision to obtain $(r^{\star},c^{\star})$, with $c^{\star}$ set to the original correct call:
$(r,\,c)\;\xRightarrow[\text{human supervision}]{\text{post–editing}}\;(r^{\star},\,c^{\star}),$ (8)
$\mathcal{L}_{\Sigma}(c^{\star})=\texttt{Success Call}.$ (9)
The finalized item is packaged as
[…]
where $D^{+}_{>2k+1}$ is the untouched suffix of $D^{+}$ (including $m^{\mathrm{final}}_{n}$). We retain $x$ only if: (i) tags/JSON are well-formed; (ii) $c^{\star}$ is executable; (iii) $r^{\star}$ correctly cites the clean–broken contrast.
3.2 Reward Design
Preliminary.
Given a model completion $C$ and a ground truth $G$, we decompose both into three (possibly empty) parts:
$C\mapsto\big(c_{\text{ref}},\;C_{\text{calls}}=\{c_{i}\}_{i=1}^{m},\;c_{\text{final}}\big),\qquad G\mapsto\big(g_{\text{ref}},\;G_{\text{calls}}=\{g_{j}\}_{j=1}^{n},\;g_{\text{final}}\big).$ (11)
Here $c_{\text{ref}}$ (reflection) is the diagnosis text wrapped in <reflect></reflect>, $C_{\text{calls}}$ is the multiset of tool calls wrapped in <call></call>s produced by the model, and $c_{\text{final}}$ is the message wrapped in <final></final>. The ground truth can alse be decomposed into three parts mentioned above.
Component scores.
We compute three component scores:
$s_{\text{ref}}=\mathrm{Sim}(c_{\text{ref}},g_{\text{ref}}),\qquad s_{\text{call}}=\mathbb{I}\!\left[\mathrm{EqualCalls}(C_{\text{calls}},G_{\text{calls}})\right],\qquad s_{\text{final}}=\mathrm{Sim}(c_{\text{final}},g_{\text{final}}),$ (12)
where $\mathrm{Sim}\!\in[0,1]$ is a semantic similarity function, and $\mathbb{I}[\cdot]$ is the indicator:
$\mathbb{I}[P]\;=\;\begin{cases}1,&\text{if $P$ is true},\\ 0,&\text{otherwise}.\end{cases}$ (13)
We say $\mathrm{EqualCalls}(C_{\text{calls}},G_{\text{calls}})$ holds iff the two sets of produced calls can be put in a one-to-one correspondence such that for every matched pair the tool name is identical and the argument is identical.
Normalization with presence masks.
Our goal is to keep the aggregated score in $[0,1]$ even when an instance specifies only a subset of targets (e.g., only <call> without <reflect> or <final>). To this end we use normalization to renormalize over the parts that actually appear in the ground truth, so the maximum remains $1$ regardless of how many parts are present.
We define
$I_{\text{r}}=\mathbb{I}[\,g_{\text{ref}}\neq\varnothing\,],\qquad I_{\text{c}}=\mathbb{I}[\,|G_{\text{calls}}|>0\,],\qquad I_{\text{f}}=\mathbb{I}[\,g_{\text{final}}\neq\varnothing\,].$ (14)
Let $(w_{\text{r}},w_{\text{c}},w_{\text{f}})\!\geq\!0$ be normalized base weights (e.g., $w_{\text{r}}+w_{\text{c}}+w_{\text{f}}=1$). We renormalize over the active parts via
$W_{\text{act}}\;=\;w_{\text{r}}I_{\text{r}}+w_{\text{c}}I_{\text{c}}+w_{\text{f}}I_{\text{f}}.$ (15)
The aggregated structure/semantics score is then
$S\;=\;\dfrac{w_{\text{r}}I_{\text{r}}\,s_{\text{ref}}\;+\;w_{\text{c}}I_{\text{c}}\,s_{\text{call}}\;+\;w_{\text{f}}I_{\text{f}}\,s_{\text{final}}}{W_{\text{act}}}.$ (16)
This normalization yields a consistent scoring standard across fully and partially supervised instances, avoiding artificial deflation of scores when some targets are absent.
Format/penalty factor.
We designed structural penalties tailored for tool-call data formats. Specifically, $P_{miss}$ accounts for cases where the tool is not invoked at all, while $P_{extra}$ and $P_{count}$ penalize redundant calls and mismatches in the total number of calls, respectively. Let
$n=|G_{\text{calls}}|,\qquad m=|C_{\text{calls}}|,$ (17)
Here $n$ and $m$ denote the number of tools invoked in the ground truth and completion calls. Define the three components:
$P_{\text{miss}}\;=\;w_{\text{ref}}\;\mathbb{I}[\,g_{\text{ref}}\!\neq\!\varnothing\wedge c_{\text{ref}}\!=\!\varnothing\,]\;+\;w_{\text{final}}\;\mathbb{I}[\,g_{\text{final}}\!\neq\!\varnothing\wedge c_{\text{final}}\!=\!\varnothing\,]\;+\;w_{\text{calls}}\;\mathbb{I}[\,n>0\wedge m=0\,],$ (18)
$P_{\text{extra}}\;=\;w_{\text{ref}}\;\mathbb{I}[\,c_{\text{ref}}\!\neq\!\varnothing\wedge g_{\text{ref}}\!=\!\varnothing\,]\;+\;w_{\text{final}}\;\mathbb{I}[\,c_{\text{final}}\!\neq\!\varnothing\wedge g_{\text{final}}\!=\!\varnothing\,]\;+\;w_{\text{calls}}\;\mathbb{I}[\,m>0\wedge n=0\,],$ (19)
$P_{\text{count}}\;=\;w_{\text{calls}}\;\mathbb{I}[\,n>0\wedge m>0\wedge n\!\neq\!m\,]\;\frac{|n-m|}{\max(n,m)}.$ (20)
Let $\mathrm{EqualCalls}$ be the schema-strict equality on bags of calls. We use a reduction factor
$r\;=\;\begin{cases}r_{\text{reduce}},&\text{if }\mathrm{EqualCalls}(C_{\text{calls}},G_{\text{calls}}),\\ 1,&\text{otherwise,}\end{cases}\qquad r_{\text{reduce}}\in(0,1].$ (21)
Aggregate the penalty as
$P_{\text{total}}\;=\;P_{\text{miss}}\;+\;\beta_{\text{extra}}\,P_{\text{extra}}\;+\;\gamma_{\text{count}}\,P_{\text{count}},$ (22)
and define the instance-wise format factor
$\mathrm{FormatFactor}(C,G)=\begin{cases}1,\text{if }P_{\text{miss}}=0\ \wedge\ P_{\text{extra}}=0\ \wedge\ P_{\text{count}}=0,\\[4.0pt] \max(0,\min(1,\,1-\lambda_{m}P_{\text{total}}r)),\text{otherwise.}\end{cases}$ (23)
Here $\beta_{\text{extra}},\gamma_{\text{count}},\lambda_{m}\!\geq\!0$ control the strength of extra-part, count-mismatch, and overall scaling penalties, respectively; $(w_{\text{ref}},w_{\text{calls}},w_{\text{final}})\!\geq\!0$ are part weights.
Core reward and backoff.
The core reward is
$R_{\text{core}}\;=\;S\cdot F.$ (24)
Early in training, $S$ contains a binary component ($s_{\text{call}}\!\in\!\{0,1\}$) and $F$ applies hard penalties; small formatting or argument errors can drive $R_{\text{core}}$ close to zero. This yields sparse or unstable gradients and large variance across samples. To stabilize learning and provide a dense shaping signal when the exact-match objective is not yet achieved, we introduce a similarity backoff:
$R_{\text{total}}\;=\;\begin{cases}\mathrm{clip}_{[0,1]}(R_{\text{core}}),&R_{\text{core}}\geq\varepsilon,\\[4.0pt] \mathrm{clip}_{[0,1]}\!\left(w_{\text{b}}\cdot\mathrm{Sim}\!\big(\mathrm{concat}(C),\,\mathrm{concat}(G)\big)\right),&\text{otherwise,}\end{cases}$ (25)
where $w_{\text{b}}\in(0,1]$ and $\mathrm{concat}(\cdot)$ linearizes the messages. We use $\mathrm{clip}_{[0,1]}(x)=\max(0,\min(1,x))$ to keep rewards bounded.
3.3 RL for Tool-Reflection-Bench
We adopt a reinforcement-learning objective for tool calling that combines two complementary ideas: (i) DAPO-style decoupled clipping [32]: we use a decoupled clipping range with different lower/upper bounds $(\varepsilon_{\text{low}},\varepsilon_{\text{high}})$ and a clip-higher policy (a looser upper bound when $r>1$ for positive advantages), and we skip uninformative prompt groups whose rollouts carry negligible learning signal; (ii) GSPO-style sequence-level importance sampling [33]: we compute the importance ratio at the sequence level and apply clipping at the same granularity as the sequence-level reward, which avoids the mismatch between token-wise importance sampling and sequence-level rewards and stabilizes optimization.
Objective.
Let $(q,a)$ denote the dialog context and the ground-truth targets, and let $\{o_{i}\}_{i=1}^{G}$ be $G$ candidates sampled from the behavior policy $\pi_{\theta_{\text{old}}}(\cdot\mid q)$. Each completion $o_{i}$ is scored by the reward in Sec. §3.2, yielding $R_{i}\in[0,1]$. We maximize a sequence-level, asymmetrically clipped objective and minimize its negative as the loss:
$\mathcal{J}_{\text{RL}}(\theta)=\mathbb{E}_{(q,a)\sim\mathcal{D},\,\{o_{i}\}\sim\pi_{\theta_{\text{old}}}(\cdot\,|\,q)}\left[\frac{1}{G}\sum_{i=1}^{G}\min\!\Big(r_{i}(\theta)\,\hat{A}_{i},\;\mathrm{clip}\!\big(r_{i}(\theta),\,1-\varepsilon_{\text{low}},\,1+\varepsilon_{\text{high}}\big)\,\hat{A}_{i}\Big)\right],$ (26)
where $\mathrm{clip}(x,a,b)=\min\{b,\max\{a,x\}\}$ and typically $\varepsilon_{\text{high}}>\varepsilon_{\text{low}}$ (“clip-higher”).
Prompt-group dynamic filtering.
DAPO skips prompt groups whose candidates provide almost no learning signal (e.g., all-correct or all-wrong). Concretely, define batch-normalized advantages and a group-level acceptance criterion:
$\hat{A}_{i}=\frac{R_{i}-\mathrm{mean}\!\big(\{R_{j}\}_{j=1}^{G}\big)}{\mathrm{std}\!\big(\{R_{j}\}_{j=1}^{G}\big)},\qquad\mathcal{S}(q,a)=\Big\{\,i\in\{1,\ldots,G\}\;:\;|\hat{A}_{i}|>\tau_{\text{adv}}\Big\},$ (27)
and require sufficient reward dispersion within the group:
$\mathrm{Var}\!\big(\{R_{i}\}_{i=1}^{G}\big)>\tau_{\text{var}}\quad\text{and}\quad 0<|\mathcal{S}(q,a)|<G.$ (28)
If (28) fails, we drop the zero-information rollouts and (optionally) draw up to $K$ additional candidates from $\pi_{\theta_{\text{old}}}$, then re-apply the filter. Only indices in $\mathcal{S}(q,a)$ contribute to the expectation in (26).
Sequence-level importance ratio.
For a completion $o_{i}=(o_{i,1},\ldots,o_{i,|o_{i}|})$, we use the geometric-mean, length-normalized importance ratio:
$r_{i}(\theta)=\left(\prod_{t=1}^{|o_{i}|}\frac{\pi_{\theta}\!\left(o_{i,t}\mid q,\,o_{i,<t}\right)}{\pi_{\theta_{\text{old}}}\!\left(o_{i,t}\mid q,\,o_{i,<t}\right)}\right)^{\!1/|o_{i}|},$ (29)
and perform clipping at the same sequence granularity as the reward (see (26)), thereby avoiding token/sequence granularity mismatch.